OpenStudy (anonymous):

need help with evaluating this integral.... (x^9)/(sqrt(3+x^5)) dx.....

OpenStudy (anonymous):

I would recommend a u substitution of u = x^5+3

OpenStudy (anonymous):

ok i started that....

OpenStudy (anonymous):

How far did you get?

OpenStudy (anonymous):

then i got (1/5)du = x^4dx ....

OpenStudy (anonymous):

Ok, though I prefer dx = du/(5x^4)

OpenStudy (anonymous):

ok.... then

OpenStudy (anonymous):

So then, what is x^5 in terms of u?

OpenStudy (anonymous):

i dont know what you mean by that

OpenStudy (anonymous):

u = x^5 + 3 so x^5 = ?

OpenStudy (anonymous):

x^5 = u - 3 right?

OpenStudy (anonymous):

u - 3

OpenStudy (anonymous):

ok. Good. Now start substituting in your integral

OpenStudy (anonymous):

so whats the x^9?... the denominator will be sqrt3?

OpenStudy (anonymous):

The denominator will be \(\sqrt{u}\)

OpenStudy (anonymous):

thats what i meant.///

OpenStudy (anonymous):

the dx will be \[\frac{1}{5x^4}du\]

OpenStudy (anonymous):

So what's left in the numerator( after you simplify)

OpenStudy (anonymous):

idk.......

OpenStudy (anonymous):

Cmon. \[\int \frac{x^9}{\sqrt{x^5+3}}dx = \int \frac{x^5 * x^4}{\sqrt{x^5+3}}dx \] \[\text{Let }u = x^5+3 \]\[\implies x^5 = u-3 \] \[\implies 5x^4dx = du \implies dx = \frac{1}{5x^4}du\] Therefore \[\int \frac{x^5 * x^4}{\sqrt{x^5+3}}dx = \int \frac{(u-3)*x^4}{\sqrt{u}} (\frac{1}{5x^4})du\]

OpenStudy (anonymous):

And we can cancel the x^4 from top and bottom. Then split the fraction into two different fractions \[\frac{u}{\sqrt{u}} - \frac{3}{\sqrt{u}}\]

OpenStudy (anonymous):

and integrate each one separately.

OpenStudy (anonymous):

what happened to the 5 in the faraction

OpenStudy (anonymous):

yes there should be a factor of 5 outside the integral , he most likely didnt mention it because any integral of a constant multiple is just the integral times that constant